4. You roll a fair six-sided dice twice and record the results, in order. The sample space is 2,6 6,6) (a) How many possible outcomes are there? (b) What is the probability that the total (sum) of the results is equal to 10? (c) Given that the total is equal to 10, what is the probability that the first roll was a 4? (d) Given that the first roll was not a 4, what is the probability that the total is equal to 10?
Answer #1

BASED ON THE GIVEN DATA AND SAMPLE SPACE, WE CAN CALCULATE

A) POSSIBLE OUTCOMES = 6*6 = 36

B) POSSIBILITIES THAT SUM OF RESULTS IS EQUAL TO 10 ARE - (6,4) , (5,5) , (4,6)

PROBABILITY THAT SUM OF RESULTS IS EQUAL TO 10 , P = 3/36 = 0.0833

C) POSSIBILITIES THAT THE FIRST ROLL WAS 4 AND TOTAL IS EQUAL TO 10 ARE - (4,6)

PROBABILITY THAT THE FIRST ROLL WAS 4 AND TOTAL IS EQUAL TO 10, P = 1/3 = 0.3333

D) POSSIBILITIES THAT THE FIRST ROLL WASN'T  4 AND TOTAL IS EQUAL TO 10 ARE - (6,4) , (5,5)

PROBABILITY THAT THE FIRST ROLL WASN'T 4 AND TOTAL IS EQUAL TO 10, P = 2/3 = 0.6667